OpenStudy (anonymous):

solve the system x/4 - y/4 = -1 x + 4y = -9

OpenStudy (zzr0ck3r):

solve the second equation for x x = -9-4y plug in -9-4y when you see an x in the first equation (1/4)(-9-4y)-y/4 = -1 (-9-4y-y) = -4 -5y=5 ....

OpenStudy (anonymous):

then you plug in -1 for your y value in the second equation?

OpenStudy (anonymous):

@zzr0ck3r